# Aligning Agents like Large Language Models

### v1

@@ Appendix A Discussion of General Procedure for Aligning Agents
We break down our procedure for aligning agents with preferences (see Figure 1) into five steps:
@@
    The first ingredient for training large language models is to train a large, scalable transformer architecture with self-supervised next-token prediction on a diverse dataset of human text, to obtain a language prior. In the context of agents on modern console games, we interpret this as imitation learning to predict the next action taken in human gameplay data, to obtain a behavioural prior. Specifically this involves training the transformer autoregressively to learn a policy $p(a_{t}|o_{t},...,o_{t-H})$. For the purpose of this work, we consider an agent trained on diverse data within a particular game, but note that given our unified observation and action spaces (visual observations and gamepad actions), it would also be possible to train across games, as explored in previous work (Reed et al., 2022; Lee et al., 2022).
@@
    The next step in the current LLM pipeline is to fine-tune the pre-trained ‘foundation’ model on task relevant data, such as instruction data (Chung et al., 2022). Pre-trained transformer models have been shown to fine-tune more effectively, essentially increasing the size of the fine-tuning data compared to training from scratch (Hernandez et al., 2021). For decision-making agents this involves fine-tuning by imitation learning on the task (or game) of interest. For specific behaviours, this could also consist of fine-tuning on demonstrations of the desired behaviour by the game designer. The training procedure is equivalent to step 1.
@@
    Fine-tuned LLMs are subsequently prompted to generate multiple responses which are compared by human labellers to provide preferences. In the context of agents, the prompt becomes the initial observation (and optionally the context of previous observations and actions). The agent is then rolled out from a given initial start state multiple times to collect multiple trajectories. Similarly to LLMs, the temperature of the softmax sampling of the policy for action selection can be increased to generate more diverse behaviours from the agent for easier comparison. A human (e.g. game designer) then provides preferences on these trajectories, such as preferring trajectories where the agent plays in a certain style.
@@
    A reward model is then trained on these online trajectories, commonly using a Bradley Terry model (Bradley & Terry, 1952) for pairwise comparisons, so that the reward model provides higher reward for preferred trajectories. While this reward model is usually trained from scratch in the context of agents, the modern LLM procedure utilises the pre-trained or fine-tuned policy model by replacing the action classification head with a scalar regression head. This enables the reward model to also benefit from the pre-training, and share the same knowledge as the agent to reduce out-of-distribution behaviours (Ziegler et al., 2020).
@@
    Finally the agent can be trained with online reinforcement learning, to maximise the reward provided by the reward model, thereby aligning the agent with the game designer’s preferences. Since online reinforcement learning can be inefficient, we find that additional preference fine-tuning can be performed on the high reward trajectories before deploying it online to get the agent closer to the desired behaviour. A common failure mode at this stage is reward model over-optimisation, where the agent performs behaviours that maximise the reward model output but are not aligned with preferences (also known as reward hacking). If this occurs, regularisation towards the original policy can be added or steps 3-5 can be repeated to generate new preferences on the reward hacking behaviour which can be used to re-train the reward model. This may take multiple iterations (e.g. 5 reward model iterations were used for Llama2 (Touvron et al., 2023)), to eventually obtain an aligned agent.
This procedure combines the benefits of large scale pre-training to obtain a widely generalisable agent, with the benefits of reinforcement learning from preferences to obtain an aligned agent. Furthermore, it also combines the benefits of offline learning to avoid the well-known sample inefficiency of online reinforcement learning, with online fine-tuning to alleviate the well-known problems associated with imitation learning agents going out of distribution (Ross et al., 2011) by enabling active online learning (Ostrovski et al., 2021).
Our procedure for aligning agents with preferences can also be summarised in algorithmic form:
@@
1:Large diverse dataset $D$, task specific dataset $T$, environment $E$, transformer policy $\pi$
2:Train policy $\pi$ on large diverse dataset $D$ with imitation learning (Pomerleau, 1991)
$\mathcal{L}_{U}(\pi)=\mathbb{E}_{D}\left[-\log\pi(a^{\tau}_{t}|o^{\tau}_{t},o^{\tau}_{t-1},...)\right]$
3:Fine-tune policy $\pi$ on task specific dataset or demonstrations $T$
$\mathcal{L}_{FT}(\pi)=\mathbb{E}_{T}\left[-\log\pi(a^{\tau}_{t}|o^{\tau}_{t},o^{\tau}_{t-1},...)\right]$
4:Run fine-tuned policy $\pi$ in environment and provide preferences on pairwise comparisons on online trajectories to obtain preference data $(\tau_{A}\succ\tau_{B})\in P$
5:Initialise additional reward model $\hat{r}$ from $\pi$ and train on preferences $P$ using Bradley-Terry model (Bradley & Terry, 1952)
@@
6:Train fine-tuned policy $\pi_{RL}$ using online reinforcement learning, using reward model $\hat{r}$ to provide return for online trajectory $\tau$ to align agent with preferences. For example, using REINFORCE (Williams, 1992):
$\mathcal{L}_{RL}(\pi_{RL})=-\mathbb{E}_{\pi_{RL}}\left[\sum_{t=1}^{T}\gamma^{t}\hat{r}(\tau)\log\pi_{RL}(a_{t}|o_{t},...)\right]$
@@ C.1 Base Model
The first layer of the conv net has kernels of shape $8\times 8$, with a stride of $4$, and a padding of $3$ and maps to $16$ channel dimension. This is followed by 4 lots of ConvNext (Liu et al., 2022) and downsampling blocks (kernel of shape $3\times 3$, stride of $2$, padding of $1$, doubling the channel dimension). Finally, a LayerNorm (Ba et al., 2016) is applied to the output.
The transformer operates on sequences of 32 timesteps using learnt positional encodings. The output of the transformer is layernormed, and then fed into an MLP with a single hidden layer of $1024$ dimensions with a GELU (Hendrycks & Gimpel, 2023) non-linearity.
For our optimiser we use AdamW (Loshchilov & Hutter, 2019) with a learning rate of $41e-$ 4 and a weight decay of $41e-$ 4. We use a batch size of 256 with a learning rate warmup period of 1000 updates and a gradient clipping value of $1$. We train with the same image augmentations as used by (Baker et al., 2022), and filter out all no-op actions.
@@ C.2 Fine-Tuning of Base Model
We train for 1500 batches of size 128 with a learning rate of $61e-$ 6 with the same image augmentations and no-op filtering as for pre-training with $200$ warmup steps.
@@ C.3 Reward Models
For the Agent Encoder model, we feed the trajectory into the fine-tuned agent and take the layernormed output of the transformer, corresponding to timesteps 00 up to $100$, as $1024$ dimensional embeddings. This is larger since it must capture the 32 context steps rather than just a single observation. The parameters of the fine-tuned agent are not trained.
@@
To train the reward models we use a minibatch of size $2048$, learning rate of $41e-$ 4, and an $L_{2}$ regularisation penalty of $0.1$. We train all models for $200$ epochs, except for the largest training set size of $1000$ trajectories which we train for $50$ epochs.
@@ C.4 Alignment Training
Preference Fine-Tuning: after training the reward model on its dataset of $M$ trajectories (which result in a dataset of up to $N=(M)(M-1)/2$ comparisions), we compute the reward for each of these $M$ trajectories. We then sort them by magnitude, and take the top $20\%$ of these as a smaller dataset to perform behaviour cloning on.
For this final step of BC, we use a learning rate of $51e-$ 5 with $1000$ updates on minibatches of size $256$. We only train the parameters of the MLP after the transformer layers.
@@
Reinforcement Learning: in our experiments we use an undiscounted REINFORCE loss on batches of 16 episodes of up to 100 timesteps. We use a learning rate of $41e-$ 4 and once again only train the parameters of the MLP after the transformer layers. If an error occurs during an episode’s rollout, we simply drop the samples from that episode and subsequently use a smaller batch size for the update.
@@ Appendix D Discussion of Offline to Online Distribution Shift
As we mention in Section 3, the nature of using a real AAA video game is such that there is significant offline to online distribution shift between our offline training data and our online evaluation. One such source is due to character selection and customisable visual modifications, as shown below in Figure 10.
@@
It is well-known that offline imitation learning often suffers from online distribution drift, particularly when learning from pixels in a continuous partially observable environment, due to accumulating errors (Ross et al., 2011). The visual distribution shift in our environment shown in Figure 10 only exacerbates these issues. As a result, imitation learning agents may not consistently perform imitated behaviours online even if the loss has converged on the offline data. This challenge is fundamental to offline learning (Ostrovski et al., 2021). This motivates the need to have some online fine-tuning (in our case from preferences) to refine the policy to consistently perform the desired behaviour online.
Figure 4 demonstrates that even after the supervised fine-tuning stage on a dataset of trajectories that always reach a jumppad and involve the same character, a significant proportion of trajectories ($\sim 10\%$) do not reach any jumppad. However, after online fine-tuning, we are able to increase the overall success rate to nearly 100% as shown in Figure 19.
Appendix E Ablation of Unsupervised Pre-Training and Model Scaling
E.1 Ablation of Unsupervised Pre-Training
We see from the results in Sections 3.1 and 3.2 that fine-tuning improves the task-specific performance of our pre-trained agent. However, to determine whether fine-tuning our base model is beneficial over simply training a model from scratch on our curated task-specific dataset, we also ablate the unsupervised pre-training stage. We train our agent from scratch for 20k updates, using the same procedure as used previously for fine-tuning (see Appendix C) until the loss appeared to converge. We subsequently evaluated the agent in the environment for 1000 episodes following the same procedure described in Section 3. The distribution of jumppads reached by the agent trained from scratch on the fine-tuning dataset is shown below (left) for comparison with the original pre-trained+fine-tuned agent (right) in Figure 11.
Figure 11: Distribution of jumppads reached by an imitation learning agent trained from scratch on the task-specific dataset used for fine-tuning the base model (left) compared to the base agent that was fine-tuned on this dataset.
We find that this agent has a greater failure rate (around 20% for the fine-tuned only agent fail to reach any jumppad compared to only around 10% for the pre-trained+fine-tuned agent) and a much narrower distribution of jumppads reached. This is surprising given the task-specific dataset consists of successful trajectories evenly split between the jumppads.
As additional anecdotal evidence for the benefit of unsupervised pre-training, we also noticed fine-tuned agents miss a jumppad, hit the wall behind and then turn around to hit the jumppad. Since the fine-tuning dataset consists of only curated trajectories that directly hit the dataset, this behaviour was not present in the fine-tuning dataset, and agents trained from scratch on the fine-tuned dataset are found to continuously run into the wall if they miss the jumppad, as they have never seen that observation before (and since moving forward is the most likely action). This behavior is shown on our project webpage at: https://adamjelley.github.io/aligning-agents-like-llms/.
This provides further anecdotal but intuitive evidence for the benefits of unsupervised pre-training.
E.2 Preliminary Model Scaling Analysis
@@
  • ($\sim$103M: $8$ layers with $1024$ hidden dims and $8$ heads for each attention layer.)
@@
The jumppad distributions for these models are provided in Figure 12. We see that these smaller models have much greater failure rates that the larger 103M parameter models shown in Figure 11, demonstrating that larger models are beneficial for imitation learning from pixels even on this relatively small task-specific dataset of 300 trajectories. While larger models still may be beneficial (particularly with pre-training on our large unsupervised dataset described in Section 3), larger models would further increase the crucial inference cost and we find that 103M parameters is sufficient for further alignment, as demonstrated in Figure 11.
E.3 Alignment of Models Trained from Scratch by Size
To complete our ablation, we now investigate how pre-training and model size affects online alignment. To do so we followed the procedure in Appendix A to generated preferences as in Section 3. We then trained corresponding reward models and compare aligning these models trained from scratch to our pre-trained and fine-tuned model, as shown below in Figure 13.
@@ E.3 Alignment of Models Trained from Scratch by Size
Overall, the results are relatively high variance due to the small dataset these results are based on. However, bearing that in mind, we can see some general trends emerging nonetheless. We see that our pre-trained model aligns significantly better than the equivalent size model trained from scratch when aligned left, but worse when aligned right (due to the lack of diversity in the pre-trained+fine-tuned model as explained in Section 3.5.2). We also see that larger models generally align better (although the bias of the 25M parameter model towards going right makes this trend less clear) since the success rate generally increases more quickly during alignment for the larger models.
We also note that in this experiment we only considered using the reward models trained on 500k comparisons, in order to see whether it was possible to successfully align a model trained solely on our small task-specific dataset. As noted in the previous section, these models have significantly less diversity in their behaviour than the model first pre-trained and then fine-tuned. This makes it much more costly (in terms of human time) to generate sufficient labelled examples of the desired behaviour to properly train the corresponding reward model (for example, less than $10\%$ of trajectories reach the left jumppad for the model trained from scratch compared to over $30\%$ for the pre-trained + fine-tuned model). This means that while these results show an upper bound in terms of reward model performance, we would expect the larger and pre-trained models to have less degradation in their alignment when using reward models trained on smaller quantities of preference data.

### v2

@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of machine learning, in particular decision-making agents. While these agents could have many positive impacts, ensuring the safety of such agents is crucial. Aligning agents with human preferences is a prominent approach to obtaining helpful and harmless agents, but this procedure has known open problems and limitations (casper_open_2023). Video games therefore provide an important test bed for such research, and we hope work such as ours can mitigate risks and provide insights in a safe environment that may generalize to other more real-world applications such as robotics.
@@ Appendix A Discussion of General Procedure for Aligning Agents
We break down our procedure for training general decision-making agents (see Figure 1) into five steps:
@@
    The first ingredient for training large language models is to train a large, scalable transformer architecture with self-supervised next-token prediction on a diverse dataset of human text, to obtain a language prior. In the context of agents on modern console games, we interpret this as imitation learning to predict the next action taken in human gameplay data, to obtain a behavioral prior. Specifically this involves training the transformer autoregressively to learn a policy $p(a_{t}|o_{t},a_{t-1}...,o_{t-H};e_{t})$ with optional text conditioning $e_{t}$ containing embedded context on the game, task or current goal. For the purpose of this work, we consider an agent trained on diverse data within a particular game, but note that given our unified observation and action spaces (visual observations and gamepad actions), it would also be possible to train across games, as explored in previous work (reed_generalist_2022; lee_multi-game_2022).
@@
    The next step in the current LLM pipeline is to fine-tune the pre-trained ‘foundation’ model on task relevant data, such as instruction data (chung_scaling_2022). Pre-trained transformer models have been shown to fine-tune more effectively, essentially increasing the size of the fine-tuning data compared to training from scratch (hernandez_scaling_2021). For decision-making agents this involves fine-tuning by imitation learning on the task (or game) of interest. For specific behaviors, this could also consist of fine-tuning on demonstrations of the desired behavior by the game designer. The training procedure is equivalent to Step 1.
@@
    Fine-tuned LLMs are subsequently prompted to generate multiple responses which are compared by human labelers to provide preferences. In the context of agents, the prompt becomes the initial observation (and optional context of previous observations, actions and/or textual instructions). The agent is then rolled out from a given initial start state multiple times to collect multiple trajectories. Similarly to LLMs, the temperature of the softmax sampling of the policy for action selection can be increased to generate more diverse behaviors from the agent for easier comparison. A human (e.g. game designer) then provides preferences on these trajectories, such as preferring trajectories where the agent plays in a certain style.
@@
    A reward model is then trained on these online trajectories, commonly using a Bradley Terry model (bradley_rank_1952) for pairwise comparisons, so that the reward model provides higher reward for preferred trajectories. While this reward model is usually trained from scratch in the context of agents, the modern LLM procedure utilizes the pre-trained or fine-tuned policy model by replacing the action classification head with a scalar regression head. This enables the reward model to also benefit from the pre-training, and share the same knowledge as the agent to reduce out-of-distribution behaviors (ziegler_fine-tuning_2020). This reward model should ideally capture general behavior preferences to provide more scalable feedback to the agent.
@@
    Finally the agent can be trained with online reinforcement learning to maximize the reward provided by the reward model, thereby aligning the agent with the game designer’s preferences. Since online reinforcement learning can be inefficient, various alternatives can be used, such as Direct Preference Optimization (DPO) (rafailov_direct_2023) with an offline preference dataset or ReST (gulcehre_reinforced_2023), but online learning is generally preferred (xu_is_2024). A common failure mode at this stage is reward model over-optimization, where the agent performs behaviors that maximize the reward model output but are not aligned with preferences (also known as reward hacking). If this occurs, regularization towards the original policy can be added or steps 3-5 can be repeated to generate new preferences on the reward hacking behavior which can be used to update the reward model. This may take multiple iterations (e.g. 5 reward model iterations were used for Llama 2 (touvron_llama_2023)), to eventually obtain an aligned agent.
This procedure combines the benefits of large scale pre-training to obtain a widely generalizable agent, with the benefits of reinforcement learning from preferences to obtain an aligned agent. Furthermore, it also combines the benefits of offline learning to avoid the renown sample inefficiency of online reinforcement learning, with the benefits of online learning to alleviate the well-known problems associated with offline agents going out-of-distribution (ross_efficient_2010; ross_reduction_2011) by enabling some active online learning (ostrovski_difficulty_2021) with a sensible behavioral prior.
Our procedure for aligning agents with preferences can also be summarized in algorithmic form:
@@
0: Large diverse dataset $D$, task specific dataset $T$, environment $E$, transformer policy $\pi$
1: Train policy $\pi$ on large diverse dataset $D$ with imitation learning (pomerleau_efficient_1991)
[…]
2: Fine-tune policy $\pi$ on task specific dataset or demonstrations $T$
[…]
3: Run fine-tuned policy $\pi$ in environment and provide preferences on pairwise comparisons on online trajectories to obtain preference data $(\tau_{A}\succ\tau_{B})\in P$
4: Initialise additional reward model $\hat{r}$ from $\pi$ and train on preferences $P$ using Bradley-Terry model (bradley_rank_1952)
@@
5: Train fine-tuned policy $\pi_{RL}$ using online reinforcement learning, using reward model $\hat{r}$ to provide return for online trajectory $\tau$ to align agent with preferences. For example, using REINFORCE (williams_simple_1992) with Kullback–Leibler divergence regularization parameterized by $\beta$:
$\mathcal{L}_{RL}(\pi_{RL})=-\mathbb{E}_{\pi_{RL}}\left[\sum_{t=1}^{T}\gamma^{t}\hat{r}(\tau)\log\pi_{RL}(a_{t}|o_{t},...)-\beta\log\left(\frac{\pi_{RL}(a_{t}|o_{t},)}{\pi_{FT}(a_{t}|o_{t},)}\right)\right]$
@@ C.1 Base Model
The first layer of the conv net has kernels of shape $8\times 8$, with a stride of $4$, and a padding of $3$ and maps to $16$ channel dimension. This is followed by 4 lots of ConvNext (liu_convnet_2022) and downsampling blocks (kernel of shape $3\times 3$, stride of $2$, padding of $1$, doubling the channel dimension). Finally, a LayerNorm (ba_layer_2016) is applied to the output.
The transformer operates on sequences of 32 timesteps using learnt positional encodings. The output of the transformer is layernormed, and then fed into an MLP with a single hidden layer of $1024$ dimensions with a GELU (hendrycks_gaussian_2023) non-linearity.
For our optimiser we use AdamW (loshchilov_decoupled_2019) with a learning rate of 1e41e-4 and a weight decay of 1e41e-4. We use a batch size of 256 with a learning rate warmup period of 1000 updates and a gradient clipping value of $1$. We train with the same image augmentations as used by (baker_video_2022), and filter out all no-op actions.
@@ C.2 Fine-Tuning of Base Model
We train for 1500 batches of size 128 with a learning rate of 1e61e-6 with the same image augmentations and no-op filtering as for pre-training with $200$ warmup steps.
@@ C.3 Reward Models
For the Agent Encoder model, we feed the trajectory into the fine-tuned agent and take the layernormed output of the transformer, corresponding to timesteps 0 up to $100$, as $1024$ dimensional embeddings. This is larger since it must capture the 32 context steps rather than just a single observation. The parameters of the fine-tuned agent are not trained.
@@
To train the reward models we use a minibatch of size $2048$, learning rate of 1e41e-4, and an $L_{2}$ regularisation penalty of $0.1$. We train all models for $200$ epochs, except for the largest training set size of $1000$ trajectories which we train for $50$ epochs.
@@ C.4 Alignment Training
Preference Fine-Tuning: after training the reward model on its dataset of $M$ trajectories (which result in a dataset of up to N=(M)(M1)/2N=(M)(M-1)/2 comparisions), we compute the reward for each of these $M$ trajectories. We then sort them by magnitude, and take the top $20\%$ of these as a smaller dataset to perform behaviour cloning on.
For this final step of BC, we use a learning rate of 1e51e-5 with $1000$ updates on minibatches of size $256$. We only train the parameters of the MLP after the transformer layers.
@@
Reinforcement Learning: in our experiments we use an undiscounted REINFORCE loss for 9600 episodes, using a batch size of 16 to give 600 parameter updates. Each episode is up to 100 timesteps or around 10s, so this corresponds to around 1 day of real time gameplay in total. We use a learning rate of 1e41e-4 and once again only train the parameters of the MLP after the transformer layers. If an error occurs during an episode’s rollout, we simply drop the samples from that episode and subsequently use a smaller batch size for the update.
@@ Appendix D Discussion of Offline to Online Distribution Shift
The nature of using a real AAA video game is such that there is significant offline to online distribution shift between our offline training data and our online evaluation, as discussed in Section 3.3. Typical visual distribution shifts in our setting, due to character selection and customizable visual modifications, are shown below in Figure 11.
@@
It is also well established that offline-only imitation learning usually suffers from online distribution drift, particularly when learning from pixels in a continuous and partially observable environment (ross_efficient_2010; ross_reduction_2011). Both of these sources of distribution drift often lead to significant accumulating error in an online trajectory. As a result, imitation learning agents may not consistently perform imitated behaviors online even if the loss has converged on the offline data. This challenge is fundamental to offline learning, as the agent cannot learn from its own observations and actions (ostrovski_difficulty_2021). This motivates the need to have some online fine-tuning (in our case from preferences) to refine the policy to consistently perform the desired behavior online.
Figure 4 demonstrates that even after supervised fine-tuning on a dataset of trajectories that always reach a jumppad and involve the same character, a significant proportion of trajectories ($\sim 10\%$) do not reach any jumppad. However, after online fine-tuning, we are able to increase the overall success rate to nearly 100% as shown in Appendix I. We note that this is analogous to recent findings for LLMs that online training (when appropriately configured) can lead to performance improvements over offline-only training (xu_is_2024).
Appendix E Preliminary Model Scaling and Alignment Analysis
E.1 Preliminary Model Scaling Analysis
@@
  • ($\sim$103M: 88 layers with $1024$ hidden dims and $8$ heads for each attention layer.)
@@
The jumppad distributions for these models are provided in Figure 12. We see that these smaller models have much greater failure rates that the larger 103M parameter models shown in Figure 5, demonstrating that larger models are beneficial for imitation learning from pixels even on this relatively small task-specific dataset of 300 trajectories. While larger models still may be beneficial (particularly with pre-training on our large unsupervised dataset described in Section 3.2), larger models would further increase the crucial inference cost and we find that 103M parameters is sufficient for further alignment, as demonstrated in Figure 5.
E.2 Preliminary Analysis of Model Alignment with Scale
To complete our ablation, we now investigate how pre-training and model size affects online alignment. To do so we followed the procedure in Appendix A to generated preferences as in Section 3.4. We then trained corresponding reward models and compare aligning these models trained from scratch to our pre-trained and fine-tuned model, as shown below in Figure 13.
@@ E.2 Preliminary Analysis of Model Alignment with Scale
Overall, the results are relatively high variance due to the small dataset these results are based on. For example, we see that our pre-trained model (brown) aligns significantly better than the equivalent size model trained from scratch (green) when aligned left, and worse when aligned right, but this is likely just due to the lack of diversity in the pre-trained+fine-tuned model as explained in Section 3.6.2. However, bearing that in mind, we can see some general trends emerging. In particular, we see some indications that larger models generally align better (ignoring the left-aligned 25M model and the right-aligned pre-trained 100M model due to the substantial left/right bias associated with these models) and allowing for differences in initial performance due to minor left/right bias, since the success rate generally increases more quickly during alignment as model size increases.
We also note that in this experiment we only considered using the original reward models trained on 500k comparisons, in order to see whether it was possible to successfully align a model trained solely on our small task-specific dataset. As noted in the previous section, these models have significantly less diversity in their behavior than the model first pre-trained and then fine-tuned. This makes it much more costly (in terms of human time) to generate sufficient labeled examples of the desired behavior to properly train the corresponding reward model (for example, less than $10\%$ of trajectories reach the left jumppad for the model trained from scratch compared to over $30\%$ for the pre-trained + fine-tuned model). This means that while these results show an upper bound in terms of reward model performance, we would expect the larger and pre-trained models to have less degradation in their alignment when using reward models trained on smaller quantities of preference data, and therefore for larger pre-trained models to perform even better in more practical settings.